OpenStudy (anonymous):

find max of function. f(x) = -x^4+2x+3. let t=x^2

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

really

OpenStudy (anonymous):

that does not help your cause at all.

OpenStudy (anonymous):

give me a medal

OpenStudy (anonymous):

maybe it would if it was \[f(x)=-x^4+2x^2+3\]

OpenStudy (anonymous):

are you sure it is just \[f(x)=-x^4+2x+3\]? because if you replace x^2 by t you get \[-t^2+2\sqrt{t}+3\] a big mess

OpenStudy (anonymous):

it says hint: let t=x^2

OpenStudy (anonymous):

can you give me a medal

OpenStudy (anonymous):

nvm i got it using calculator

OpenStudy (anonymous):

-2t-1/radt=0 2t=1/radt t^3/2=1/2 t=(1/2)^2/3 x=((1/2)^2/3)^2 ARE YOU CRITICAL POINTS AND NOW TAKE 2ND DER to find if its a max or min

OpenStudy (anonymous):

x=((1/2)^2/3)^1/2

OpenStudy (anonymous):

@bnut056 apparently this is not a calc course

OpenStudy (anonymous):

ha opps then i dont know what hes asking